# Learning in Structured Stackelberg Games

### v1

@@ 4.1 Learning without Contexts
Leveraging results from the data-driven algorithm design literature ([4, 8, 7]), such as Lemma 3.10 in [6] and Proposition 4.3 we have the following result:
@@
The pseudo-dimension of the hypothesis class, $\pazocal{P}$ satisfies $\textnormal{Pdim}(\pazocal{P})\leq 9|\pazocal{A}|\log\left(4|\pazocal{A}||\pazocal{A}_{f}|^{2}\right)$.
So we also get the generalization guarantee from Pollard [33] and Dudley [18] with function:
@@
$\epsilon_{\pazocal{G}}(m,\delta)=O\left(U\sqrt{\frac{|\pazocal{A}|\log(|\pazocal{A}||\pazocal{A}_{f}|^{2})}{m}}+U\sqrt{\frac{\ln(1/\delta)}{m}}\right),$
where $U$ is the maximum payoff value for the leader.
Discussion.
We note that the above generalization result is particularly strong. Informally, it asserts that no matter what the distribution over the follower types is, if the leader has so far encountered a number of followers that scales linearly with the size of her action set and logarithmically with the follower’s action set, it suffices for her to play a strategy that maximizes her average utility over the followers seen so far. This result holds even for an infinite number of follower types and extends prior work on learning in Stackelberg games without contextual information [26].
4.2 Learning with Contexts
We now shift to analyzing the setting where the leader has access to a data set consisting of tuples of contexts and follower types, $S_{m}=\{(\mathbf{z}_{1},f_{1}),\ldots,(\mathbf{z}_{m},f_{m})\}$. Given new contextual information, the leader aims to choose a strategy that maximizes her payoff against a follower, whose type is unknown. Here we introduce the study of the statistical complexity of distributional learning in Stackelberg games with side information.
Consider a distribution $\mathscr{D}$ over $\pazocal{Z}\times\mathscr{K}$. We define the leader’s payoff as a function of a policy $\pi:\pazocal{Z}\to\Delta(\pazocal{A})$ with respect to $\mathscr{D}$ as:
and the leader’s average payoff on a sample $S_{m}=\{(\mathbf{z}_{1},f_{1}),\ldots,(\mathbf{z}_{m},f_{m})\}\in(\pazocal{Z}\times\mathscr{K})^{m}$ as:
The goal of the learner is to deploy an algorithm, $\textsc{Alg}:\cup_{m=0}^{\infty}(\pazocal{Z}\times\mathscr{K})^{m}\to\Delta(A)^{\pazocal{Z}}$, such that given a training sequence $S_{m}=\{(\mathbf{z}_{1},f_{1}),\cdots,(\mathbf{z}_{m},f_{m})\}$, the algorithm outputs a policy $\pi_{\textsc{Alg}}:\pazocal{Z}\to\Delta(A)$ that maximizes $u_{\mathscr{D}}(\pi)$.
Suppose that the leader has access to a set of hypotheses, $\pazocal{H}$, such that each hypothesis in $\pazocal{H}$, maps contexts to follower types ($h:\pazocal{Z}\to\mathscr{K}$). We define the true error function of a hypothesis $h$ with respect to a distribution, $\mathscr{D}$, as:
Similarly, we define the empirical error of a hypothesis $h$ on a sample $S_{m}$ as:
$\text{Err}(h)=\frac{1}{m}|\{i\in[m]:h(\mathbf{z}_{i})\neq f_{i}\}|$
and the approximation error of a hypothesis class $\pazocal{H}$ on a distribution $\mathscr{D}$ to be:
Definition 4.6 (Natarajan Dimension).
Let $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$ be a hypothesis class and let $S\subseteq\pazocal{Z}$. We say that $\pazocal{H}$ N-shatters $S$ if there exist $\phi_{1},\phi_{2}:S\to\mathscr{K}$ such that $\forall\mathbf{z}\in\pazocal{Z}\;\phi_{1}(\mathbf{z})\neq\phi_{2}(\mathbf{z})$ and for any $T\subseteq S$ there exists a $g\in\pazocal{H}$ such that:
We define the realizable sample complexity of a Stackelberg-learning algorithm Alg for a hypothesis class $\pazocal{H}$ to be the function $m_{\textsc{Alg},\pazocal{H}}^{r}$ such that for every $\epsilon,\delta>0$, $m_{\textsc{Alg},\pazocal{H}}^{r}$ is the minimal integer such that for every $m\geq m_{\textsc{Alg},\pazocal{H}}^{r}$ and every distribution $\mathscr{D}$ over $\pazocal{Z}$, we have:
4.2.1 Vanilla ERM learner
We define an Empirical Risk Minimization (ERM) learner for the hypothesis class $\pazocal{H}$ as an algorithm, such that for any sample $S_{m}$ returns a function that minimizes the empirical error of that function relative to any other function in $\pazocal{H}$. Formally, we say:
Proposition 4.7.
Consider an algorithm $\textsc{Alg}_{\text{ERM}}:\cup_{n=0}^{\infty}(\pazocal{Z}\times\mathscr{K})^{n}\to\Delta(\pazocal{A})^{\pazocal{Z}}$ that outputs the policy:
where $h_{ERM}$ satisfies equation (3). Then for any such $h_{\text{ERM}}$, there exists constant $C>0$, such that for every $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$,
We present the proof of Proposition 4.7 in Appendix B.
4.2.2 Stackelberg ERM learner
In this section, we show how to improve upon the sample complexity bound shown in Proposition 4.7, by leveraging the interaction between the Stackelberg game and the hypothesis class. Consider the following hypothesis class of possible leader policies:
$\Pi_{\pazocal{H},\pazocal{G}}=\Bigl{\{}\pi_{h}:\pazocal{Z}\to\Delta(\pazocal{A})|h\in\pazocal{H},\pi_{h}(\mathbf{z})=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x}))\Bigl{\}}$
and the following variation of an ERM rule.
Definition 4.8 (Stackelberg ERM).
Consider a hypothesis class $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$ and a Stackelberg game, $\pazocal{G}$. We define the Stackelberg ERM learning algorithm, $\textsc{Alg}_{\text{S-ERM}}:\cup_{n=0}^{\infty}(\pazocal{Z}\times\mathscr{K})^{n}\to\Delta(\pazocal{A})^{\pazocal{Z}}$, such that given a training sequence $S_{m}=\{(\mathbf{z}_{1},f_{1}),\cdots,(\mathbf{z}_{m},f_{m})\}$, outputs the policy $\pi_{\text{S-ERM}}:\pazocal{Z}^{m}\to\Delta(\pazocal{A})$ that satisfies:
$\pi_{\text{S-ERM}}\in\arg\min_{\pi\in\Pi_{\pazocal{H},\pazocal{G}}}\frac{1}{m}\sum_{i=1}^{m}\mathbbm{1}\{\pi(\mathbf{z}_{i})\neq\mathbf{x}_{i}^{*}\},$ (4)
where $\mathbf{x}_{i}^{*}=\arg\max_{x\in\Delta(\pazocal{A})}\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{i},\mathbf{x},b_{f_{i}}(\mathbf{z}_{i},\mathbf{x}))$.
Definition 4.9 (Stackelberg-Natarajan Dimension).

### v2

@@ 4.1 Learning without Contexts
Using Lemma˜4.3 and results from the data-driven algorithm design literature [6, 12, 14], we immediately get the following bound on the pseudo-dimension.
Theorem 4.5 (Adaptation of Lemma 3.10 in Balcan et al. [9]).
$\textnormal{Pdim}(\pazocal{P})\leq 9|\pazocal{A}|\log\left(4|\pazocal{A}||\pazocal{A}_{f}|^{2}\right)$.
It is known that the finiteness of the pseudo-dimension of a function class is a sufficient condition to show a generalization guarantee for the class [34, 24]. In our case, the above bound implies a generalization guarantee with function
$\epsilon_{\pazocal{G}}(m,\delta)=O\left(\sqrt{\frac{|\pazocal{A}|\log(|\pazocal{A}||\pazocal{A}_{f}|^{2})}{m}}+\sqrt{\frac{\ln(1/\delta)}{m}}\right).$
Discussion.
We note that the above generalization result is particularly strong. Informally, it asserts that for any distribution over the follower types, if the leader has so far encountered a number of followers that scales linearly with the size of her action set and logarithmically with the follower’s action set, it suffices for her to play a strategy that maximizes her average utility over the followers seen so far. This result holds even for an infinite number of follower types and extends prior work on learning in Stackelberg games without contextual information [29].
4.2 Learning with Contexts: Problem Setting and Definitions
In this section, we introduce the study of the statistical complexity of distributional learning in structured Stackelberg games. The leader has access to a data set consisting of tuples of contexts and follower types drawn from a fixed and unknown distribution. Given new contextual information drawn from the same distribution, the leader aims to choose a strategy that maximizes her payoff against a follower whose type is unknown. As in the online setting, this comes with the promise that there exists a function, $h^{*}$, in a class, $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$, that perfectly predicts each follower type given the context. We denote a learning algorithm for distributional learning in structured Stackelberg games as a mapping $A:(\pazocal{Z}\times[K])^{*}\to\Delta(\pazocal{A}){Z}$. Given a context, $\mathbf{z}$, we write the leader’s loss with respect to $h^{*}$, when playing strategy $\hat{\mathbf{x}}\in\Delta(\pazocal{A})$ as
$r(\mathbf{z},\hat{\mathbf{x}},f^{(h^{*}(\mathbf{z}))})=\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u\left(\mathbf{z},\mathbf{x},b_{f^{(h^{*}(\mathbf{z}))}}(\mathbf{z},\mathbf{x})\right)-u\left(\mathbf{z},\hat{\mathbf{x}},b_{f^{(h^{*}(\mathbf{z}))}}(\mathbf{z},\hat{\mathbf{x}})\right)$
We handle two variants of the notion of sample complexity, which will be central in showing our results. We begin with the classic PAC sample complexity, which we adapt to our setting.
Definition 4.6 (PAC Sample complexity).
The PAC sample complexity $m(\pazocal{H},\pazocal{G};\epsilon,\delta)$ of a function class $\pazocal{H}$ with respect to a Stackelberg game $\pazocal{G}$ is defined as $m(\pazocal{H},\pazocal{G};\epsilon,\delta)=\inf_{A}m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$, where the infimum is over all possible learning algorithms. With $m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$ we denote the minimal integer such that for any $m\geq m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$, every distribution $\mathscr{D}$ on $\pazocal{Z}$, and any true target hypothesis $h^{*}\in\pazocal{H}$, the expected loss of $A$,
$\mathop{\mathbb{E}}_{\mathbf{z}\sim\mathscr{D}}\left[r(\mathbf{z},A(S)(\mathbf{z}),f^{(h^{*}(\mathbf{z}))})\right],$
is at most $\epsilon$ with probability $1-\delta$ over the set $S=\{(\mathbf{z}_{i},f^{(h^{*}(\mathbf{z}_{i}))})\}_{i\in[m]}$, where $\forall i\in[m]\;\mathbf{z}_{i}\sim\mathscr{D}$.
Definition 4.7 (PAC Cut-off sample complexity).
The PAC sample complexity $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$ of a function class $\pazocal{H}$ with respect to a Stackelberg game $\pazocal{G}$ is defined as $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)=\inf_{A}m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$, where $m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$ is the minimal integer such that for any $m\geq m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$, every distribution $\mathscr{D}$ on $\pazocal{Z}$, and any true target hypothesis $h^{*}\in\pazocal{H}$, the expected cut-off loss of algorithm $A$,
$\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}[r(\mathbf{z},A(S)(\mathbf{z}),f^{(h^{*}(\mathbf{z}))})>\gamma],$
is at most $\epsilon$ with probability $1-\delta$ over the set $S=\{(\mathbf{z}_{i},f^{(h^{*}(\mathbf{z}_{i}))})\}_{i\in[m]}$, where $\forall i\in[m]\;\mathbf{z}_{i}\sim\mathscr{D}$.
For the remainder of this section our results will be with respect to the PAC cut-off sample complexity. Any bound with respect to this benchmark immediately implies a bound on the PAC sample complexity as well. It is not hard to generalize the result of Lemma˜4.8 to our setting, so we omit the proof.
Lemma 4.8 (Lemma 1 in [3]).
For every $\epsilon,\delta\in(0,1)^{2}$ and every $\pazocal{H}\subseteq[0,1]^{\pazocal{Z}}$, it holds that
$m(\pazocal{H};\sqrt{\epsilon},\delta,\sqrt{\epsilon})\leq m(\pazocal{H};\epsilon,\delta)\leq m(\pazocal{H},\epsilon/2,\delta,\epsilon/2).$
4.3 Lower bound
In this section, we introduce the $\gamma$-valued Stackelberg-Natarajan dimension, a new measure of complexity that characterizes a structured Stackelberg game and, as we will see, controls the lower bound of the sample complexity. Indeed, we prove that the finiteness of this measure is a necessary condition for learnability in our setting. This definition is inspired by the classic Natarajan dimension from multiclass prediction [31, 21], and more recently, the $\gamma$-Natarajan dimension from realizable regression [3]. We start with introducing the notion of a $\gamma$-SN-shattered set. Informally, a set of contexts is $\gamma$-SN-shattered if $\pazocal{H}$ can fit complex patterns of follower types on the set. Further, these follower types must be sufficiently different, such that if the type is chosen adversarially among them, the leader incurs loss at least $\gamma$ regardless of her strategy.
Definition 4.9 ($\gamma$-SN-Shattered Set).
Let $\pazocal{H}:[K]^{\pazocal{Z}}$ and $\pazocal{G}$ be a Stackelberg game. We say that a set $S\in\pazocal{Z}^{n}$ is $\gamma$-SN-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$ if there exist $g_{0},g_{1}:\pazocal{Z}\to[K]$ such that
$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{g_{0}(\mathbf{z}_{i}),g_{1}(\mathbf{z}_{i})\}}r(\mathbf{z}_{i},\mathbf{x},f^{(j)})>\gamma\;\;\forall\mathbf{z}_{i}\in S$
and for all $b\in\{0,1\}^{n}$ there exists an $h_{b}\in\pazocal{H}$ such that $h_{b}(\mathbf{z}_{i})=\begin{cases}g_{0}(\mathbf{z}_{i})\;\;\text{if}\;\;b_{i}=0\\ g_{1}(\mathbf{z}_{i})\;\;\text{if}\;\;b_{i}=1\end{cases}$.
Definition 4.10 ($\gamma$-valued Stackelberg-Natarajan Dimension).
The $\gamma$-valued Stackelberg-Natarajan dimension, denoted as $\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$, is the cardinality of the largest set that is $\gamma$-SN-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$.
Before stating our main result of this section, we point out that it is important that the two functions, $g_{0}$ and $g_{1}$ be $\gamma$-far from each other in a minmax sense. This allows us to prove a lower bound that holds for any learning algorithm.
@@ 4.3 Lower bound
Let $A$ be any learning algorithm and $\epsilon,\delta,\gamma\in(0,1)^{3}$ such that $\delta<\epsilon$, then
@@
$m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)\geq\Omega\left(\frac{\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})+\log(1/\delta)}{\epsilon}\right).$
Proof.
Let $n_{0}=\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$. Then we know that for any $1\leq n\leq n_{0}$ there exists a set $S_{n}\in\pazocal{Z}^{n}$ and functions $g_{0},g_{1}:\pazocal{Z}\to\Delta(\pazocal{A})$ such that for any set $T\subseteq S$ there exists an $h$ that agrees with $g_{0}$ on all $\mathbf{z}\in T$ and with $g_{1}$ for any $\mathbf{z}\in S\setminus T$. We will show that for a given learning algorithm $A:(\pazocal{Z}\times[K])^{*}\to\Delta(\pazocal{A}){Z}$, there exists a distribution $\mathscr{D}$ on $\pazocal{Z}$ such that the algorithm requires at least $C_{0}\frac{\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})+\log(\frac{1}{\delta})}{\epsilon}$ samples. To do that, we will construct a distribution with support the $n_{0}$-sized $\gamma$-shattered set $S$. We take $\delta<1/15$ and define the distribution $\mathscr{D}$ on $S$ as follows.
$\mathbb{P}[\mathbf{z}_{1}]=1-16\epsilon,\;\;\;\;\mathbb{P}[\mathbf{z}_{i}]=\frac{16\epsilon}{d-1}\;\forall i\in\{2,\dots,d-1\}.$
We draw a set $X$ of $m=\frac{d-1}{64\epsilon}$ i.i.d. samples from this distribution. We let $A$ be an algorithm that takes $X$ as the input and then produces a mapping $A(X):\pazocal{Z}\to\Delta(\pazocal{A})$. We define
$\text{er}_{\mathscr{D},\gamma,h}(A(X)):=\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}[r(\mathbf{z},A(X)(\mathbf{z}),f^{(h(\mathbf{z}))})>\gamma].$
It suffices to establish that there exists $h^{*}\in\pazocal{H}$ such that $\mathbb{P}_{X\sim\mathscr{D}^{m}}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]>1/15$. First consider the event
$B:|\{\mathbf{x}\in X:\mathbf{x}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}\}|\leq(d-1)/2.$
Using Markov’s inequality, it is easy to show that $\mathbb{P}[B]\geq 1/2$. Let $U=\{\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}\setminus X$ be the points excluding $\mathbf{z}_{1}$ that are unseen by the algorithm through the training set. Given $B$ we know that $|U|\geq\frac{d-1}{2}$. For the remainder of this section we will focus on
$\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X)):=\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}[r(\mathbf{z},A(X)(\mathbf{z}),f^{(h(\mathbf{z}))})>\gamma\land\mathbf{z}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n_{0}}\}],$
which trivially lower bounds $\text{er}_{\mathscr{D},\gamma,h}(A(X))$. Also, we consider the following way to choose a true hypothesis $h$. First, we choose vector $b$ uniformly at random from the set $\{0,1\}^{n_{0}}$. We then set $h(\mathbf{z}_{i})=g_{0}(\mathbf{z}_{i})$ for all $b_{i}=0$ and $h(\mathbf{z}_{i})=g_{1}(\mathbf{z}_{i})$ for all $b_{i}=1$. We note that choosing a random $h$ in this way is equivalent to flipping a fair coin for each point in $S$ to determine its label. Since for any point $\mathbf{z}_{i}\in U$ both labels $g_{0}(\mathbf{z}_{i})$ and $g_{1}(\mathbf{z}_{i})$ are consistent with the training set $X$ (by definition of an SN-shattered set), we have that $A(X)$ is independent of the labeling of points in $U$. We also know that if we condition on the point drawn from $\mathscr{D}$ being some unseen point $\mathbf{u}\in U$, we get that for any strategy $A(X)(\mathbf{u})\in\Delta(\pazocal{A})$ of the learner
$\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\geq\begin{cases}1\;\;\text{if}\;h(\mathbf{u})=\arg\max_{j\in\{g_{0}(\mathbf{u}),g_{1}(\mathbf{u})\}}r(\mathbf{u},A(X)(\mathbf{u}),f^{(j)})\\ 0\;\;\text{otherwise}.\end{cases}$
Since we are randomly choosing $h(\mathbf{z})$ to be $g_{0}(\mathbf{z}_{i})$ or $g_{1}(\mathbf{z}_{i})$, we have that the two conditions above happen with probability 1/21/21 / 2 each. Given that there are at least $(d-1)/2$ unseen points, each sampled with probability $\frac{16\epsilon}{d-1}$, and each forcing the learner to make a $\gamma$-error with probability 1/21/21 / 2, we get that
$\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\;|\;B\right]>\frac{1}{2}\cdot\frac{16\epsilon}{d-1}\cdot\frac{(d-1)}{2}=4\epsilon.$
From this analysis and the probability of $B$, we get that
$\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\right]$ $=\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\;|\;B\right]\cdot\mathbb{P}[B]$
$>4\epsilon\cdot\frac{1}{2}=2\epsilon.$
So there must exists some $h^{*}$ such that $\mathop{\mathbb{E}}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))\right]>2\epsilon$. We take $h^{*}$ as the target concept and show that $A$ is likely to produce high error rate. We know that
$\mathbb{P}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]\cdot\mathop{\mathbb{E}}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))\;|\;\text{er}_{\mathscr{D},\gamma,h}(A(X))>\epsilon\right]$
$+(1-\mathbb{P}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right])\cdot\mathop{\mathbb{E}}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))\;|\;\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))\leq\epsilon\right]>2\epsilon$
Given that for any algorithm, $A$, it holds that $\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\leq\mathbb{P}[\mathbf{z}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}]=16\epsilon$ we get
$\mathop{\mathbb{E}}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\;|\;\text{er}_{\mathscr{D},\gamma,h}(A(X))>\epsilon\right]\leq 16\epsilon\;\;\text{for any}\;\;h$
which implies that for $h^{*}$
$\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]\cdot 16\epsilon+(1-\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right])\cdot\epsilon>2\epsilon$
$\iff\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]>1/15.$
Therefore, the learner needs at least $m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)\geq\frac{d-1}{64\epsilon}$ samples.
We now consider the case, where $X$ contains only $\mathbf{z}_{1}$. This occurs with probability $(1-16\epsilon)^{m}\geq e^{-32m\epsilon}$, which is greater than $\delta$, when $m\leq\frac{\log(1/\delta)}{32\epsilon}$. Therefore we get that for any algorithm $A$:
$m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)\geq\max\biggl{\{}\frac{d-1}{64\epsilon},\frac{\log(1/\delta)}{16\epsilon}\biggr{\}}=C_{0}\frac{\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})+\log(\frac{1}{\delta})}{\epsilon}$
Lastly, we consider the case, where $\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})=\infty$. Then for any shattered set of size $n\in\mathbb{N}$, we can repeat the argument above, which would show that the cut-off sample complexity of any algorithm is at least $\Omega\left(\frac{n+\log(\frac{1}{\delta})}{\epsilon}\right)$, thus establishing that $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)=\infty$. ∎
4.4 Upper bound
In this section, we introduce the $\gamma$-valued Stackelberg-Graph dimension, which, as we will see, is a sufficient condition for distributional learning in structured Stackelberg games. It is also a natural relaxation of the $\gamma$-Natarajan dimension and an adaptation of similar dimensions from classification [31, 21] and regression [3] to our setting.
Definition 4.12 ($\gamma$-SG-Shattered Set).